Adriana has $15.00 in dimes and quarters.
a. Name 3 ordered pairs (d,q) of dimes and quarters she might have.
b. Do these points lie on the same line?

Answer by htmentor(1343)

Adriana has $15.00 in dimes and quarters.
a. Name 3 ordered pairs (d,q) of dimes and quarters she might have.
b. Do these points lie on the same line?
=======================
a.
If q=10, d=(1500-25*10)/10=125 (125,10)
If q=30, d=(1500-25*30)/10=75 (75,30)
If q=50, d=(1500-25*50)/10=25 (25,50)
b. Yes. We can write an equation for the line.
Since the total amount is 1500, and dimes are worth 10 and quarters are worth 25, we have
10d + 25q = 1500
Solve for q:
q = (1500-10d)/25 = 60 - 0.4d